Prove that c2=a2+b2- 2abcosy is equal to a= bcosy+√c2- b2×sin2y
1
Expert's answer
2019-08-05T12:04:23-0400

We have quadratic equation


"c^2=a^2+b^2-2abcos(y)"

"a^2-a(2bcos(y))+b^2-c^2=0"

Using quadratic formula


"a_{1,2}=bcos(y)±\sqrt{b^2cos^2(y)-(b^2-c^2)}"

"a_{1,2}=bcos(y)±\sqrt{b^2cos^2(y)-b^2+c^2}"

"a_{1,2}=bcos(y)±\sqrt{-b^2(1-cos^2(y))+c^2}"

"a_{1,2}=bcos(y)±\sqrt{-b^2sin^2(y)+c^2}"


"a_{1,2}=bcos(y)±\sqrt{c^2-b^2sin^2(y)}"

So this is not strict equivalence:

"if\ a=bcos(y)+\sqrt{c^2-b^2sin^2(y)},\"

then a is the root of the equation

"c^2=a^2+b^2-2abcos(y)"


if a is the root of the equation


"c^2=a^2+b^2-2abcos(y)"

then


"a=bcos(y)+\sqrt{c^2-b^2sin^2(y)}"

or


"a=bcos(y)-\sqrt{c^2-b^2sin^2(y)}."